# Mathematical Language Models: A Survey

### 14 Dec 2023 v2

@@ 2.2.2 Theorem Proving
Theorem proving (TP) refers to the process of demonstrating that a statement is correct based on existing axioms and facts, which is a long-term challenge in AI [10.1145/1455567.1455605]. In the context of LLMs, people began to try to use the huge knowledge base of LLMs and some manual guidance to solve this task. LLMs like GPT-3 [brown2020language] can process natural text to understand the premises of the theorem, and apply certain logical reasoning rules to complete the proof requirements. This can then be done by calling relevant tools or using manual inspection to check the correctness of the process. The advanced approach aims to leverage LLM’s superior computing power and vast knowledge base relative to humans to automate the theorem-proving process, a task that previously required professional mathematicians.
@@
In the field of AI, research on theorem proving has progressed from data sets containing large numbers of human-written proofs (such as CoqGym [yang2019learning]) to complex models that can autonomously generate proof strategies (such as ASTactic [yang2019learning]). This improvement is reflected in the application of some language models that are based on Transformer [vaswani2017attention] to TP tasks, especially in recent research such as GPT-f [polu2020generative]. Most notably, some of the proofs generated by these models have been formally recognized by the mathematical community. Jiang et al. [jiang2022draft] further refined this process by using informal proofs to guide TP models. NaturalProofs [welleck2021naturalproofs] extends these capabilities by leveraging the language of natural mathematics to create a rich corpus of model training and evaluation. At the same time, DeepMath [irving2016deepmath] and INT [wu2020int] have promoted the development of this field by demonstrating the effectiveness of neural sequence models for premise selection and evaluating the LLMs’ generalization ability in theorem proving.
@@
Ongoing research is delving into the integration of language models and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction [DBLP:journals/corr/abs-2009-03393, han2021proof, lample2022hypertree, jiang2022thor], automating formalization processes [DBLP:conf/nips/WuJLRSJS22, jiang2022draft], and developing unified tools [llmstep]. Due to the substantial computational demands of exploration, language models applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models [first2023baldur, jiang2022draft]. LLEMMA [azerbayev8llemma] demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
@@
The above series of progressive research work shows that utilizing the powerful computing power and extensive knowledge base of LLMs is the main development direction for future TP tasks. The more excellent models (such as GPT-4 [openai2023gpt4]) obviously have better reasoning capabilities without any human intervention, but we also need to be clear about one thing, that is the inevitable hallucination problem in neural text generation models [maynez-etal-2020-faithfulness]. The hallucination problem refers to the generated content that is nonsensical or unfaithful to the provided source content [ji2023survey]. The occurrence of hallucinations at any step in the reasoning process will lead to errors in the final reasoning results. Therefore, solving the hallucination problem is also particularly important in future work related to TP.
3 PLMs-based Methods
@@
Based on the transformer architecture [vaswani2017attention] with self-attention mechanisms, the “pre-training and fine-tuning" paradigm has revolutionized the field of natural language processing. Within the context of Pre-trained Language Models (PLMs), numerous approaches have been proposed to tackle text generation problems [noorbakhsh2021pretrained]. Autoregressive LMs (ALMs, e.g., GPT-1 [radford_gpt-1_2018] and T5 [rael_exploring_nodate]) and Non-Autoregressive LMs (NALMs, e.g., BERT [kenton2019bert] and Roberta [liuRoBERTaRobustlyOptimized2019]) are the two primary strategies for mathematics (Figure 2).
@@
LMs  (§3.1) [ GPT-$f$ [polu2020generative], EPT [kim_point_2020] , Generate &\&& Rank [shen_generate_2021] , Thor [jiang2022thor] , HTPS [lample_hypertree_nodate] , Galactica [taylor_galactica_2022] , MATH-PLM [hendrycks2021measuring] , LISA [jiang2021lisa]
PACT [han2021proof] , Minerva [lewkowycz_solving_nodate] , LIME [wu_lime_nodate] , leaf, text width=54.2em ] ] [ Non-Autoregression
LMs  (§3.2) [ Aristo [clarkRegentsScienceExams2021], GenBERT [gevaInjectingNumericalReasoning2020] , NF-NSM [fengInjectingNumericalReasoning2022] , MWP-BERT [liangMWPBERTNumeracyaugmentedPretraining2022] , MathBERT[pengMathBERTPretrainedModel2021] , TAGOP [zhu2021tat] , MT2Net [zhao2022multihiertt]
DeductReasoner [jieLearningReasonDeductively2022] , BERT-TD [liSeekingPatternsNot2022] , leaf, text width=54.2em ] ] ] [ Large
@@
Building [ Auto-explanation [nakamoto2023enhancing], ProofNet [azerbayev2023proofnet] , WizardLM [xu_wizardlm_2023] , Wizardmath [luo2023wizardmath] , leaf, text width=44.6em ] ] [ Instruct
Tuning [ MathGLM [yang2023gpt], Goat [liu_goat_2023], Calculon [muffo_evaluating_nodate], PaLM 2-L-Math [liu_improving_2023], LLEMMA [azerbayev8llemma] , leaf, text width=44.6em ] ] [ In-context
Learning [ ScratchpadGPT [nye_show_2021] , Codex-math [drori2022neural] , In-Context Sampling [liu2022makes] , ComplexPrompt [fu2022complexity]
ART [zhang2023automatic] , leaf, text width=44.6em ] ] ] [ Tool-based
@@
Methods [ SymbLLM [he2023solving], PoT [chen2022program], Codex-math [drori2022neural] , MathPrompter [imani2023mathprompter] , PAL [gao2023pal] , Code-scratchpad [upadhyay-etal-2023-improving] , leaf, text width=44.6em ] ] [ Multi-tool
Methods [ Toolformer [schick2023toolformer], Chameleon [lu2023chameleon] , ART [paranjape2023art] , ToolkenGPT [hao2023toolkengpt] , CRITIC [gou2023critic] , TALM [parisi2022talm]
Tool-Documentation [hsieh2023tool] , leaf, text width=44.6em ] ] ] [ Fundamental CoT
@@
CoT [ Answer Rationales [ling2017program] , MathQA [amini2019mathqa], CoT [wei2022chain], Complexity-based CoT [fu2022complexity] , LVSS [lightman2023lets], RFT [yuan2023scaling]
Thought Propagation [yu2023thought] , AFT [wang2023making] , PoT [chen2022program] , AOT [sel2023algorithm] , MAmmoTH [yue2023mammoth] , leaf, text width=44.6em ] ] [ Construction of
CoT [ Zero-shot CoT [kojima_large_nodate] , Auto-CoT [zhang2023automatic] , Complexity-based CoT [fu2022complexity] , PromptPG-CoT [lu2022dynamic]
AutoMate CoT [shum_automatic_2023] , BoostedPrompt [pitis_boosted_2023] , leaf, text width=44.6em ] ] ] [ Advanced CoT
@@
Methods [ Code-based self-verification [zhou2023solving] , VerifyCoT [ling2023deductive] , DIVERSE [li2023making] , Verify-and-Edit [zhao2023verify]
@@
Retrieval-CoT [he_rethinking_2022] , SCREWS [shridhar2023screws] , leaf, text width=44.6em ] ] [ Ensemble-based
@@
Methods [ Self-Consistency [wang2022self] , Diversity-of-Thought [naik2023diversity] , Complexity-based CoT [fu2022complexity], DIVERSE [li2023making]
@@
Self-check [miao2023selfcheck], MCR [yoran2023answering], Rank-verifier [cobbe2021training], GRACE [khalifa2023discriminator] , leaf, text width=44.6em ] ] [ Planing-based
Methods [ ToT [yao2023tree] , TouT [mo2023tree] , ToT-system [long2023large] , GoT [yao2023beyond, besta_graph_2023, lei2023boosting] , RESPROMPT [jiang2023resprompt]
RAP [hao2023reasoning] , LATS [zhou2023language] , LLM+P [liu2023llm+] , LLM+DP [dagan2023dynamic] , Self-refine [madaan2023self, sun2023adaplanner] , ISR-LLM [zhou2023isr]
Reflexion [shinn2023reflexion] , leaf, text width=44.6em ] ] [ Socratic Teaching
@@
Methods [ Socratic Models [zeng2022socratic], SOCRATIC QUESTIONING [qi2023art], Socratic prompt [chang2023prompting], SocratiQ [al2023socratic]
multi-turn Socratic advice [ang2023socratic] , leaf, text width=44.6em ] ] ] ] ]
Figure 2: Taxonomy of language models for mathematic.
Figure 3: The distinctions between Autoregression LMs and Non-Autoregression LMs.
3.1 Autoregression LMs
@@
The distinctions between Autoregression LMs (ALMs) and Non-Autoregressive LMs (NALMs) in sequence generation tasks are depicted in Figure 3. Furthermore, ALMs can be categorized into two architectures: causal decoder (e.g., GPT-1 [radford_gpt-1_2018] and GPT-2 [radford_gpt-2_2019]) and encoder-decoder (e.g., T5 [rael_exploring_nodate]). The generation mechanisms of these two structures for sequences are depicted on the left and in the middle of Figure 3, respectively. These two architectures have not only significantly propelled the advancement of PLMs but have also become the predominant foundational architectures for LLMs.
@@
To enhance the performance of PLMs in mathematical problems, many improved methods based on ALMs have been proposed. GPT-$f$ [polu2020generative] presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. [kim_point_2020] introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank [shen_generate_2021] introduces a novel ranking task for MWPs, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor [jiang2022thor] introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS [lample_hypertree_nodate] presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica [taylor_galactica_2022] proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU [hendryckstest2021] and MATH [hendrycks2021measuring] benchmarks.
@@
Furthermore, MATH-PLM [hendrycks2021measuring] introduces a new dataset named MATH, consisting of challenging competition mathematics problems. It is found that relying solely on increasing budgets and model parameter counts would be impractical for achieving robust mathematical reasoning. LISA [jiang2021lisa] gathers an extensive collection of lemmas and theorems extracted from the Isabelle standard library and the Archive of Formal Proofs (AFP). Using this vast corpus, it builds LMs that prove theorems effectively within the AFP. PACT [han2021proof] extracts abundant self-supervised data from kernel-level proof terms, enabling joint training alongside the standard tactic prediction objective. This effectively enhances the model’s ability for TP. Minerva [lewkowycz_solving_nodate] trains a language model, initially pre-trained on general natural language data, through additional training on technical content to address Quantitative Reasoning Problems. It achieves a SOTA performance on technical benchmarks including MATH [hendrycks2021measuring]. LIME [wu_lime_nodate] introduces a novel pre-training methodology that specifically learns inductive bias for mathematical reasoning. Observations indicate that models trained using LIME exhibit superior performance compared to standard transformers.
3.2 Non-Autoregression LMs
@@
Unlike ALMs, NALMs enable the model to generate all parts of a sequence simultaneously, without depending on previously generated content, as depicted on the right side of Figure 3. Notably, models such as BERT [kenton2019bert], Roberta [liuRoBERTaRobustlyOptimized2019], and related architectures utilize masked word representations, leveraging pre-trained context-aware embeddings as comprehensive semantic features. This significantly elevates the performance benchmarks of NLP tasks.
@@
In the field of mathematics, researchers have proposed various methods for designing models to address mathematical reasoning and computational problems. For example, Aristo [clarkRegentsScienceExams2021] fine-tunes BERT using scientific curriculum data, yielding promising results on science exams. GenBERT [gevaInjectingNumericalReasoning2020] and NF-NSM [fengInjectingNumericalReasoning2022] enhance the numerical reasoning capabilities of models by incorporating numerical data into the training process of PLMs. MWP-BERT [liangMWPBERTNumeracyaugmentedPretraining2022] further enhances the model’s capacity to represent and calculate numerical values by incorporating numeric attributes into symbol placeholders. MathBERT[pengMathBERTPretrainedModel2021] employs additional joint training of text and formulas to effectively capture the semantic-level structural information of formulas. TAGOP [zhu2021tat], MT2Net [zhao2022multihiertt], and DeductReasoner [jieLearningReasonDeductively2022] utilize BERT or RoBERTa to extract the fundamental arithmetic relationships between quantities, enabling mathematical reasoning and operations. BERT-TD [liSeekingPatternsNot2022] utilizes semantic encoding and contrastive learning to cluster problems with similar prototype equations, thereby enhancing the understanding of MWP patterns.
4 LLMs-based Methods

### 23 Feb 2024 v3

@@ 2.2.2. Theorem Proving
Theorem proving (TP) refers to the process of demonstrating that a statement is correct based on existing axioms and facts, which is a long-term challenge in AI (Newell et al., 1957). In the context of LLMs, people began to try to use the huge knowledge base of LLMs and some manual guidance to solve this task. LLMs like GPT-3 (Brown et al., 2020) can process natural text to understand the premises of the theorem, and apply certain logical reasoning rules to complete the proof requirements. This can then be done by calling relevant tools or using manual inspection to check the correctness of the process. The advanced approach aims to leverage LLM’s superior computing power and vast knowledge base relative to humans to automate the theorem-proving process, a task that previously required professional mathematicians.
@@
In the field of AI, research on theorem proving has progressed from data sets containing large numbers of human-written proofs (such as CoqGym (Yang and Deng, 2019)) to complex models that can autonomously generate proof strategies (such as ASTactic (Yang and Deng, 2019)). This improvement is reflected in the application of some language models that are based on Transformer (Vaswani et al., 2017) to TP tasks, especially in recent research such as GPT-f (Polu and Sutskever, 2020a). Most notably, some of the proofs generated by these models have been formally recognized by the mathematical community. Jiang et al. (Jiang et al., 2023b) further refined this process by using informal proofs to guide TP models. NaturalProofs (Welleck et al., 2021) extends these capabilities by leveraging the language of natural mathematics to create a rich corpus of model training and evaluation. At the same time, DeepMath (Irving et al., 2016) and INT (Wu et al., 2021) have promoted the development of this field by demonstrating the effectiveness of neural sequence models for premise selection and evaluating the LLMs’ generalization ability in theorem proving.
@@
Ongoing research is delving into the integration of language models and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction (Polu and Sutskever, 2020b; Han et al., 2022; Lample et al., 2022; Jiang et al., 2022), automating formalization processes (Wu et al., 2022; Jiang et al., 2023b), and developing unified tools (Welleck and Saha, 2023). Due to the substantial computational demands of exploration, language models applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models (First et al., 2023; Jiang et al., 2023b). LLEMMA (Azerbayev et al., [n. d.]) demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
@@
The above series of progressive research work shows that utilizing the powerful computing power and extensive knowledge base of LLMs is the main development direction for future TP tasks. The more excellent models (such as GPT-4 (OpenAI, 2023)) obviously have better reasoning capabilities without any human intervention, but we also need to be clear about one thing, that is the inevitable hallucination problem in neural text generation models (Maynez et al., 2020). The hallucination problem refers to the generated content that is nonsensical or unfaithful to the provided source content (Ji et al., 2023). The occurrence of hallucinations at any step in the reasoning process will lead to errors in the final reasoning results. Therefore, solving the hallucination problem is also particularly important in future work related to TP.
3. PLMs-based Methods
@@
Based on the transformer architecture (Vaswani et al., 2017) with self-attention mechanisms, the “pre-training and fine-tuning” paradigm has revolutionized the field of natural language processing. Within the context of Pre-trained Language Models (PLMs), numerous approaches have been proposed to tackle text generation problems (Noorbakhsh et al., 2021). Autoregressive LMs (ALMs, e.g., GPT-1 (Radford et al., 2018) and T5 (Raffel et al., [n. d.])) and Non-Autoregressive LMs (NALMs, e.g., BERT (Kenton and Toutanova, 2019) and Roberta (Liu et al., 2019)) are the two primary strategies for mathematics (Figure 2).
@@
LMs  (§3.1) [ GPT-$f$ (Polu and Sutskever, 2020a), EPT (Kim et al., 2020) , Generate &\&& Rank (Shen et al., 2021) , Thor (Jiang et al., 2022) , HTPS (Lample et al., [n. d.]) , Galactica (Taylor et al., 2022) , MATH-PLM (Hendrycks et al., 2021b) , LISA (Jiang et al., 2021)
PACT (Han et al., 2022) , Minerva (Lewkowycz et al., [n. d.]) , LIME (Wu et al., [n. d.]) , leaf, text width=54.2em ] ] [ Non-Autoregression
LMs  (§3.2) [ Aristo (Clark et al., 2021), GenBERT (Geva et al., 2020) , NF-NSM (Feng et al., 2022) , MWP-BERT (Liang et al., 2022) , MathBERT(Peng et al., 2021) , TAGOP (Zhu et al., 2021) , MT2Net (Zhao et al., 2022)
DeductReasoner (Jie et al., 2022) , BERT-TD (Li et al., 2022) , leaf, text width=54.2em ] ] ] [ Large
@@
Building [ Auto-explanation (NAKAMOTO et al., 2023), ProofNet (Azerbayev et al., 2023) , WizardLM (Xu et al., 2023) , Wizardmath (Luo et al., 2023) , leaf, text width=44.6em ] ] [ Instruct
Tuning [ MathGLM (Yang et al., 2023), Goat (Liu and Low, 2023), Calculon (Muffo et al., [n. d.]), PaLM 2-L-Math (Liu et al., 2023b), LLEMMA (Azerbayev et al., [n. d.]) , leaf, text width=44.6em ] ] [ In-context
Learning [ ScratchpadGPT (Nye et al., 2021) , Codex-math (Drori et al., 2022) , In-Context Sampling (Liu et al., 2022) , ComplexPrompt (Fu et al., 2023)
ART (Zhang et al., 2023c) , leaf, text width=44.6em ] ] ] [ Tool-based
@@
Methods [ SymbLLM (He-Yueya et al., 2023), PoT (Chen et al., 2022a), Codex-math (Drori et al., 2022) , MathPrompter (Imani et al., 2023) , PAL (Gao et al., 2023) , Code-scratchpad (Upadhyay et al., 2023) , leaf, text width=44.6em ] ] [ Multi-tool
Methods [ Toolformer (Schick et al., 2023), Chameleon (Lu et al., 2023a) , ART (Paranjape et al., 2023) , ToolkenGPT (Hao et al., 2023b) , CRITIC (Gou et al., 2023a) , TALM (Parisi et al., 2022)
Tool-Documentation (Hsieh et al., 2023) , leaf, text width=44.6em ] ] ] [ Fundamental CoT
@@
CoT [ Answer Rationales (Ling et al., 2017) , MathQA (Amini et al., 2019), CoT (Wei et al., 2022), Complexity-based CoT (Fu et al., 2023) , LVSS (Lightman et al., 2023), RFT (Yuan et al., 2023a)
Thought Propagation (Yu et al., 2023a) , AFT (Wang et al., 2023a) , PoT (Chen et al., 2022a) , AOT (Sel et al., 2023) , MAmmoTH (Yue et al., 2023) , leaf, text width=44.6em ] ] [ Construction of
CoT [ Zero-shot CoT (Kojima et al., [n. d.]) , Auto-CoT (Zhang et al., 2023c) , Complexity-based CoT (Fu et al., 2023) , PromptPG-CoT (Lu et al., 2023b)
AutoMate CoT (Shum et al., 2023) , BoostedPrompt (Pitis et al., 2023) , leaf, text width=44.6em ] ] ] [ Advanced CoT
@@
Methods [ Code-based self-verification (Zhou et al., 2023d) , VerifyCoT (Ling et al., 2023) , DIVERSE (Li et al., 2023b) , Verify-and-Edit (Zhao et al., 2023a)
@@
Retrieval-CoT (He et al., 2022) , SCREWS (Shridhar et al., 2023) , leaf, text width=44.6em ] ] [ Ensemble-based
@@
Methods [ Self-Consistency (Wang et al., 2023b) , Diversity-of-Thought (Naik et al., 2023) , Complexity-based CoT (Fu et al., 2023), DIVERSE (Li et al., 2023b)
@@
Self-check (Miao et al., 2023), MCR (Yoran et al., 2023), Rank-verifier (Cobbe et al., 2021), GRACE (Khalifa et al., 2023) , leaf, text width=44.6em ] ] [ Planing-based
Methods [ ToT (Yao et al., 2023b) , TouT (Mo and Xin, 2023) , ToT-system (Long, 2023) , GoT (Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023) , RESPROMPT (Jiang et al., 2023a)
RAP (Hao et al., 2023a) , LATS (Zhou et al., 2023e) , LLM+P (Liu et al., 2023a) , LLM+DP (Dagan et al., 2023) , Self-refine (Madaan et al., 2023; Sun et al., 2023) , ISR-LLM (Zhou et al., 2023b)
Reflexion (Shinn et al., 2023) , leaf, text width=44.6em ] ] [ Socratic Teaching
@@
Methods [ Socratic Models (Zeng et al., 2022), SOCRATIC QUESTIONING (Qi et al., 2023), Socratic prompt (Chang, 2023), SocratiQ (Al-Hossami et al., 2023b)
multi-turn Socratic advice (Ang et al., 2023) , leaf, text width=44.6em ] ] ] ] ]
Figure 2. Taxonomy of language models for mathematic.
Figure 3. The distinctions between Autoregression LMs and Non-Autoregression LMs.
3.1. Autoregression LMs
@@
The distinctions between Autoregression LMs (ALMs) and Non-Autoregressive LMs (NALMs) in sequence generation tasks are depicted in Figure 3. Furthermore, ALMs can be categorized into two architectures: causal decoder (e.g., GPT-1 (Radford et al., 2018) and GPT-2 (Radford et al., 2019)) and encoder-decoder (e.g., T5 (Raffel et al., [n. d.])). The generation mechanisms of these two structures for sequences are depicted on the left and in the middle of Figure 3, respectively. These two architectures have not only significantly propelled the advancement of PLMs but have also become the predominant foundational architectures for LLMs.
@@
To enhance the performance of PLMs in mathematical problems, many improved methods based on ALMs have been proposed. GPT-$f$ (Polu and Sutskever, 2020a) presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. (Kim et al., 2020) introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank (Shen et al., 2021) introduces a novel ranking task for MWPs, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor (Jiang et al., 2022) introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS (Lample et al., [n. d.]) presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica (Taylor et al., 2022) proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) benchmarks.
@@
Furthermore, MATH-PLM (Hendrycks et al., 2021b) introduces a new dataset named MATH, consisting of challenging competition mathematics problems. It is found that relying solely on increasing budgets and model parameter counts would be impractical for achieving robust mathematical reasoning. LISA (Jiang et al., 2021) gathers an extensive collection of lemmas and theorems extracted from the Isabelle standard library and the Archive of Formal Proofs (AFP). Using this vast corpus, it builds LMs that prove theorems effectively within the AFP. PACT (Han et al., 2022) extracts abundant self-supervised data from kernel-level proof terms, enabling joint training alongside the standard tactic prediction objective. This effectively enhances the model’s ability for TP. Minerva (Lewkowycz et al., [n. d.]) trains a language model, initially pre-trained on general natural language data, through additional training on technical content to address Quantitative Reasoning Problems. It achieves a SOTA performance on technical benchmarks including MATH (Hendrycks et al., 2021b). LIME (Wu et al., [n. d.]) introduces a novel pre-training methodology that specifically learns inductive bias for mathematical reasoning. Observations indicate that models trained using LIME exhibit superior performance compared to standard transformers.
3.2. Non-Autoregression LMs
@@
Unlike ALMs, NALMs enable the model to generate all parts of a sequence simultaneously, without depending on previously generated content, as depicted on the right side of Figure 3. Notably, models such as BERT (Kenton and Toutanova, 2019), Roberta (Liu et al., 2019), and related architectures utilize masked word representations, leveraging pre-trained context-aware embeddings as comprehensive semantic features. This significantly elevates the performance benchmarks of NLP tasks.
@@
In the field of mathematics, researchers have proposed various methods for designing models to address mathematical reasoning and computational problems. For example, Aristo (Clark et al., 2021) fine-tunes BERT using scientific curriculum data, yielding promising results on science exams. GenBERT (Geva et al., 2020) and NF-NSM (Feng et al., 2022) enhance the numerical reasoning capabilities of models by incorporating numerical data into the training process of PLMs. MWP-BERT (Liang et al., 2022) further enhances the model’s capacity to represent and calculate numerical values by incorporating numeric attributes into symbol placeholders. MathBERT(Peng et al., 2021) employs additional joint training of text and formulas to effectively capture the semantic-level structural information of formulas. TAGOP (Zhu et al., 2021), MT2Net (Zhao et al., 2022), and DeductReasoner (Jie et al., 2022) utilize BERT or RoBERTa to extract the fundamental arithmetic relationships between quantities, enabling mathematical reasoning and operations. BERT-TD (Li et al., 2022) utilizes semantic encoding and contrastive learning to cluster problems with similar prototype equations, thereby enhancing the understanding of MWP patterns.
4. LLMs-based Methods